OpenStudy (anonymous):

True or false questionsssss The solution to -16 = x = -30 is -14 true or false The solution to m = (-38) = 15 is 53 True or false The solution to -35x = -105 is -3 True or false The solution to w/-7 = -21 is 3 True or false last but not least The solution to b - (-27) = -30 is -57 True or False?

OpenStudy (anonymous):

@thisSucks could you help at all? :)

OpenStudy (anonymous):

or @John_ES

OpenStudy (john_es):

What means, "The solution to -16 = x = -30 is -14"? In particular, what means -16=x=-30, are there two equals in the original problem?

OpenStudy (anonymous):

Oh! I'm sorry it was +* not = I forgot to hit shift! Lol it's that better?

OpenStudy (john_es):

Then you mean, -16 + x = -30 m + (-38) = 1?

OpenStudy (john_es):

Could you rewrite them with the correct symbols, it would be helpful.

OpenStudy (anonymous):

The very first question is "The solution to -16 + x = -30 is -14 True or False" that's what it says

OpenStudy (anonymous):

the other ones are written correctly I'm pretty sure :)

OpenStudy (john_es):

Ok, then, 1) True, -16-14=-30 2) m+(-38)=15, m=15+38, m=53, True 3) x= -105/(-35)=3, so it is false, 4)False, because (check the equation is ok) \[\frac{w}{-7}=-21\Rightarrow w=84\] 5) b-(-27)=-30, b=-30-27=-57, True

OpenStudy (anonymous):

Thank you so much! I don't know why math is sooooo hard for me, it just is.

OpenStudy (john_es):

Keep the hard work, the trick is doing a LOT of problems ;), I can assure you it.

OpenStudy (anonymous):

Awe, Thank you I will :)

OpenStudy (anonymous):

U already got your answers I think